OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

y-y1 = m(x-x1) <- right ?

ganeshie8 (ganeshie8):

We have have point (4, -2) and slope -3

ganeshie8 (ganeshie8):

so, lets start by writin the equation in point-slope form, after that we can canvert it to standard form

OpenStudy (anonymous):

i believe so. I got this x + 3y = −10 but i dont think its correct

ganeshie8 (ganeshie8):

lets see, equation in point slope form :- y-y1 = m(x-x1)

ganeshie8 (ganeshie8):

(4, -2) and slope -3 x1 y1 m

ganeshie8 (ganeshie8):

y --2 = -3(x-4) y+2 = -3x+12 3x+y = 10

OpenStudy (anonymous):

So its 3x+y=10

ganeshie8 (ganeshie8):

Yes ! you must have flipped sign somehwere... thats all i guess

OpenStudy (anonymous):

thankyou, now this one ( its a hw assighment and its worth 100 points) I wanna make sure I get them correct. Given that f(x) = 6x + 2 and g(x) = \[\frac{ 2x+4 }{ 5 }\] solve for g(f(1)).

OpenStudy (anonymous):

for this one I got 3

ganeshie8 (ganeshie8):

wrong, try again

OpenStudy (anonymous):

ok brb

ganeshie8 (ganeshie8):

you wanto find : g(f(1)) first find f(1)

ganeshie8 (ganeshie8):

okie take ur time

OpenStudy (anonymous):

f is 8 correct?

ganeshie8 (ganeshie8):

you mean f(1) is 8 ? then you're correct :)

ganeshie8 (ganeshie8):

g(f(1)) g(8)

ganeshie8 (ganeshie8):

now find g(8)

OpenStudy (anonymous):

yes! 1 is 8 ! ok im gonna find g(8)

ganeshie8 (ganeshie8):

good luck !

OpenStudy (anonymous):

its 1 ! :o

ganeshie8 (ganeshie8):

no try again

OpenStudy (anonymous):

what ! -.-

OpenStudy (anonymous):

ok ok let me write it on paper that helps

ganeshie8 (ganeshie8):

\(\large g(x) = \frac{ 2x+4 }{ 5 } \)

OpenStudy (anonymous):

I got it ! it's 4 g=4

ganeshie8 (ganeshie8):

\(\large g(8) = \frac{ 2(8)+4 }{ 5 } \)

ganeshie8 (ganeshie8):

\(\large g(8) = \frac{16+4 }{ 5 } \)

ganeshie8 (ganeshie8):

yup ! sorry my computer was behaving bit..

OpenStudy (anonymous):

so its not 4?

ganeshie8 (ganeshie8):

g(f(1)) = g(8) = 4

OpenStudy (anonymous):

@ganeshie8 sorry i took so long i had a teacher calling (online schooling)

OpenStudy (anonymous):

Now this one What is the equation of a horizontal line passing through the point (2, 10)? I got x=2

ganeshie8 (ganeshie8):

equaiton of horizontal line :- y = k

ganeshie8 (ganeshie8):

since its going thru (2, 10) equation wud be y = 10

ganeshie8 (ganeshie8):